OpenStudy (anonymous):

what does the notation 3.41E-7 mean?

OpenStudy (anonymous):

E is 10^

OpenStudy (anonymous):

so any number can follow after the carrot symbol which indicates an exponent

OpenStudy (anonymous):

so for examples 2E2 = 2* 10^2 = 200

OpenStudy (anonymous):

because 10^2 = 100 and multiplying by 2 you get 200

OpenStudy (anonymous):

this is called scientific notation so 3.41E-7 is 3.14 * 10^-7 or 0.000000314

OpenStudy (anonymous):

a decimal can be written like that form too example would be a negative exponent multiplied by number equal to or great than 1 2E-2 = 2* 10^-2 = 0.02 or can be written 2 / 10^2